Louisiana State University Health Sciences Center Administration & Finance
 

Thursday, November 16, 2006

  • The manufacturer of the Dental School (temporary) chillers attempted to start up the chillers today.  Neither would start and the technician spent a considerable amount of time checking the electrical wiring because both chillers seem to have the same problem. 
    • Two parts (one for each chiller) have been ordered.  One should be delivered on Monday, and the other should arrive later in the week.
    • Loop pump #1 was aligned today and is now within tolerances.  
    • Condenser water Pump #2 has bad packing and leaks.  Repairs will be made as required. 
  • The project to install new lighting in the Student Parking Lots continued. 
    • In Student Lot #2, the new electrical panel was installed.  Work also continued to install the steel reinforcement “cages” and the concrete forms around each of the timber pilings in preparation for pouring the concrete light pole bases this weekend.
    • In Student Lot #3, the area around the entrance/exit was cleaned up.  All holes in the entire lot have been backfilled.  The new light poles are scheduled to be installed on Monday or Tuesday of next week.
  • At 1542 Tulane Ave. , the mechanical sub-contractor is still repairing pipe leaks and replacing nipples in the basement.
    • The contractor is cleaning out the vault under the Lakeside lot to make space for a future electrical riser. 
  • The emergency generator automatic transfer switch at Stanislaus Hall was repaired and successfully tested today.  It is ready for inspection by the State Fire Marshall.